OpenStudy (anonymous):

It takes 5 seconds for sound to travel 1 mile through air. It takes only 1 second for sound to travel 1 mile through water. How many minutes faster will a sound 90 miles away travel through water than air? a. 4 minutes c. 6 minutes b. 5 minutes d. 7 minutes

OpenStudy (dj3strella):

To find how many minutes sooner it would arrive, calculate the 2 times. d = r*t t = d/r In water, r = 1 mi/sec In air, r = 1/5 mi/sec 90 miles in air at 0.2 mi/sec --> 90/0.2 = 450 seconds

OpenStudy (dj3strella):

so its D